# Goal-Oriented Communications for Real-time Inference with Two-Way Delay

### v1

@@ IV-A Co-design Problem Formulation
The learning and communications co-design problem with time-variable packet length selection is formulated as:
@@
Problem (14) is an infinite-horizon average-cost SMDP, which is more challenging to solve compared to problem (6). We formulate a Bellman optimality equation for this SMDP and then simplify it using a structural result regarding the optimal policy. The time complexity for solving the simplified Bellman optimality equation using dynamic programming algorithms is substantially lower than that for solving the original Bellman optimality equation.
@@ IV-B Optimal Solution to (14)
Each ACK reception time $A_{i}$ is a decision time for the SMDP (14). The system state at time slot $A_{i}$ is represented by a tuple $\big{(}\delta,d,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$, $d=d(A_{i})$, and $c_{i}$ is the delay state in the $i$-th epoch. The actions to be taken at each decision time $A_{i}$ are the waiting time $\tau_{i+1}$, the packet length $l_{i+1}$, and the buffer position $b_{i+1}$. A detailed description of the SMDP can be found in Appendix C. The Bellman optimality equation of this SMDP can be expressed as:
@@
where $h_{2}\big{(}\delta,d,c_{i}\big{)}$ is the relative value function.
By solving the Bellman optimality equation (IV-B) using dynamic programming, we can obtain a solution to problem (14). The time complexity of this dynamic programming algorithm is high, as it requires the joint optimization of three variables for each state $\big{(}\delta,d,c_{i}\big{)}$. Theorem 2 provides a threshold rule for determining the waiting time action at each decision time $A_{i}$ and simplifies the Bellman optimality equation. The time complexity for solving the simplified Bellman optimality equation is significantly lower than that for solving (IV-B).
@@
where $\gamma\big{(}\cdot\big{)}$ is defined in (9). Furthermore, the optimal packet length $l_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ are determined by solving the simplified Bellman optimality equation
@@
Theorem 2 is proved motivated by the observation that the optimal waiting time $\tau_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ for any state $\big{(}\delta,d,c_{i}\big{)}$ can be determined by solving separate optimization problems, given the optimal packet length $l_{i+1}^{*}$. Subsequently, the solution to the waiting time problem is derived as an index-based threshold rule, which allows us to simplify the Bellman optimality equation (IV-B) to (2). The detailed proof is provided in Appendix E. ∎
Both Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [20, Chapter 11.4.4]. Policy iteration is efficient for solving SMDPs [17, 20]. The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B). The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2). The policy evaluation step is identical for both equations.
@@ V Simulations
To evaluate the performance of the optimal scheduling policies outlined in Theorems 1 and 2, we consider two experiments: (i) remote inference of autoregressive (AR) processes, which provides a model-based evaluation; and (ii) cart-pole state prediction, which provides a trace-driven evaluation.
@@ V-A Model-Based Evaluation
Control systems with memory and delay [43, 44, 45] and wireless channels [46] are often modeled as higher-order autoregressive (AR) linear time-invariant systems. Motivated by this, we consider a remote inference problem for an AR process, where the target evolves according to:
@@
where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$ and $Y_{t}\in\mathbb{R}$. Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of the target $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$. We set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$. The remaining coefficients are zero. We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$. In this experiment, the goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$. Since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic, the optimal inference error is achieved by a linear MMSE estimator. Fig. 3 shows the average inference error for AoI values $\delta=1,2,\ldots,40$ and packet lengths $l=1,2,4,6,8,10$.
Figure 3: Inference error of the AR process experiment for AoI values $\delta=1,2,\ldots,40$ and packet lengths $l=1,2,4,6,8,10$.
Using this AR process experiment, we illustrate the performance of the optimal policy derived for the scheduling problem (6)-(7) under time-invariant packet length selection. Let the number of delay states $N$ be $2$. The transmission delay $T_{i}(l)$ corresponding to each state is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1\\ \lceil 5\sigma l\rceil,&\text{if }c_{i}=2\end{cases},$ (19)
where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$. We adjust the delay memory using the parameter $\alpha=p_{12}+p_{21}$, which is equal to the sum of transition probabilities of the Markov chain according to which the delay state evolves. We set $p_{12}=p_{21}$, ensuring that the fraction of epochs with convenient delay conditions is $\frac{1}{2}$ in the steady state for any $\alpha$. The maximum allowed packet length is 10, and the buffer size $B=75$.
@@
    Policy outlined in Theorem 1 for $l=1$.
@@
The result in Fig. 4 highlights that the first policy achieves a time-average inference error up to 6 times smaller compared to the third policy. Although part of this gain can be obtained with the packet length $l=1$ by optimally determining the waiting times and buffer positions as in the second policy, a larger portion of the gain comes from selecting the appropriate packet length, accounting for the interplay between the transmission delay statistics and the packet length $l$.
@@
    Policy outlined in Theorem 1 for $l=5$.
@@
    Policy outlined in [17, Theorem 1] for a fixed packet length of $l=5$. During the simulations, this policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
The result in Fig. (5) emphasizes the importance of considering delay memory when designing scheduling algorithms. It demonstrates that the first policy offers minimal improvement compared to the second policy when $\alpha$ is close to $1$. This is because the delay distribution becomes exactly IID at $\alpha=1$. However, as $\alpha$ deviates from this value, the delay memory increases, and the advantage of the first policy becomes apparent, with a significant performance gain of up to $11.6\%$.
@@ V-B Trace-Driven Evaluation
We consider the OpenAI CartPole-v1 task [47], where a Deep Q-Network (DQN) reinforcement learning algorithm [48] is used to control the force applied to a cart to prevent the attached pole from falling over. A pre-trained DQN neural network from [5] is employed for this control task. A time-series dataset containing the pole angle $\psi_{t}$ and cart position $Z_{t}$ is obtained from [5]. This dataset was generated by simulating 10,000 episodes of the OpenAI CartPole-v1 environment using the DQN controller. The objective is to predict the pole angle and position $Y_{t}=(\psi_{t},Z_{t})$ at time $t$ based on the feature vector $X_{t-\delta}=\big{(}(\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta+l-1},Z_{t-\delta+l-1})\big{)}$, which represents a sequence of past pole angles and cart positions generated $\delta$ time slots ago, with a sequence length of $l$.
Figure 6: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and feature vector lengths $l=1,2,4,6,8,10$.
@@
Ensuring the safe operation of industrial and robotic systems often requires evaluating their state based on observations from remote sources. Consequently, the prediction problem is critical for monitoring the safety of these systems. We focus on a specific scenario where a camera remotely monitors a cart-pole system. The camera captures video frames and transmits them to a receiver. Due to the high dimensionality of the video data, transmitting the sequence of frames requires multiple time slots. Consequently, the most recently delivered sequence of video frames is generated $\delta$ time slots ago. Once received, the frames undergo signal processing to extract relevant information about the cart’s position and the pole’s angle. We assume this signal processing is accurate. Then, a predictor takes the most recent available observations and predicts the current pole angle and cart position $Y_{t}=(\psi_{t},Z_{t})$.
@@
The predictor in this experiment is a Long Short-Term Memory (LSTM) neural network consisting of one input layer, one hidden layer with 64 LSTM cells, and a fully connected output layer. The first 72% of the dataset is used for training, and the remaining data is used for inference. The performance of the trained LSTM predictor is evaluated on the inference dataset. The inference error is shown in Fig. 6 for AoI values $\delta=1,2,\ldots,80$ and feature vector lengths $l=1,2,4,6,8,10$. We observe that the inference error is not a monotonic function of AoI $\delta$ for a given feature vector length $l$. Additionally, the inference error decreases as the feature vector length $l$ increases for a fixed AoI value $\delta$.
@@
Using this cart-pole experiment, we illustrate the performance of the optimal policy outlined in Theorem 2, which is derived for the scheduling problem (14) under time-variable packet length selection. We again consider two delay states, i.e., ${N=2}$. The transmission delay $T_{i}(l)$ is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2\end{cases},$ (20)
@@
where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is identical to that in the AR process experiment. We set the parameter $\alpha=p_{12}+p_{21}=\frac{1}{20}$. The transition probabilities $p_{12}$ and $p_{21}$ are equal, ensuring that the fraction of epochs with convenient delay conditions is $\frac{1}{2}$ in steady state for any $\alpha$.
@@ VI Conclusion
This paper studies a goal-oriented communication design problem for remote inference, where an intelligent model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location. We derive two optimal scheduling policies under time-invariant and time-variable packet length selection. These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory. Finally, through the simulations, we demonstrate the significant benefits of adopting such a goal-oriented communication strategy for remote inference.
@@ Appendix B Proof of Theorem 1
By some algebraic manipulations, we can obtain

### v3

@@ IV-A Co-design Problem Formulation
The learning and communications co-design problem with general, time-variable packet length selection is formulated as
@@
Problem (14) is an infinite-horizon average-cost SMDP, which is more challenging to solve compared to problem (6). We will first formulate a Bellman optimality equation for this SMDP and then simplify it using a structural result regarding the optimal policy. The time complexity for solving the simplified Bellman optimality equation using dynamic programming algorithms is substantially lower than that for solving the original Bellman optimality equation.
@@ IV-B Optimal Solution to (14)
In the SMDP (14), the system state at each decision time $A_{i}$ is given by the tuple $\big{(}\Delta(A_{i}),d(A_{i}),C_{i}\big{)}$, where $\Delta(A_{i})$ denotes the AoI at time slot $A_{i}$, $d(A_{i})$ indicates the length of the most recently received packet by that time, and $C_{i}$ represents the delay state in the $i$-th packet transmission epoch. A realization of the system state is denoted by $\big{(}\delta,d,c_{i}\big{)}$. The actions taken at each decision time $A_{i}$ include the waiting time $\tau_{i+1}$, the packet length $l_{i+1}$, and the buffer position $b_{i+1}$. A detailed description of this SMDP can be found in Appendix C. The Bellman optimality equation for the SMDP (14) can be expressed as
@@
where $\delta\in\mathbb{Z}^{+}$, $d\in\{1,2,\ldots,B\}$, $c_{i}\in\{1,2,\ldots,N\}$, and $h_{2}\big{(}\delta,d,c_{i}\big{)}$ is the relative value function.
By solving the Bellman optimality equation (IV-B) using dynamic programming, we can obtain a solution to (14). The time complexity of this dynamic programming algorithm is high, as it requires the joint optimization of three variables for each state $\big{(}\delta,d,c_{i}\big{)}$. The following Theorem 2 provides a threshold rule for determining the waiting time action at each $A_{i}$, simplifying the Bellman optimality equation.
@@
where $\gamma\big{(}\cdot\big{)}$ is defined in (9). Furthermore, the optimal packet length $l_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ are determined by solving the following simplified Bellman optimality equation:
@@
Theorem 2 is proved based on the observation that the optimal waiting time $\tau_{i+1}^{*}$ and buffer position $b_{i+1}^{*}$ for any state $\big{(}\delta,d,c_{i}\big{)}$ can be determined by solving separate optimization problems, provided that the optimal packet length $l_{i+1}^{*}$ is known. We derive a solution for the separate waiting time optimization problem as an index-based threshold rule, which, in turn, enables the simplification of the Bellman optimality equation (IV-B) to (2). The detailed proof is presented in Appendix E. ∎
The Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [23, Chapter 11.4.4]. In particular, policy iteration is efficient for solving SMDPs [17, 23]. The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B): The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2). The policy evaluation step is identical for both equations.
@@ V Simulations
To evaluate the performance of the optimal scheduling policies in Theorems 1 and 2, we consider two experiments: (i) remote inference of auto-regressive (AR) processes, which offers a model-based evaluation; and (ii) cart-pole state prediction, which provides a trace-driven evaluation.
@@ V-A Model-Based Evaluation
Figure 3: Inference error of the AR process for AoI values ${\delta=1,2,\ldots,50}$ and packet lengths $l=1,2,4,6,8,10$.
Higher-order auto-regressive (AR) linear time-invariant systems are widely used models for various practical scenarios, such as control systems incorporating memory and delay [43, 44, 45], and wireless communication channels [46]. Motivated by this, we consider a remote inference problem for an AR process, where the target signal evolves according to:
@@
where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$ and $Y_{t}\in\mathbb{R}$. Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of the target $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$. We set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$. The remaining coefficients of the AR($10$) process are zero. We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$. In this experiment, the goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$. A linear MMSE estimator achieves the optimal inference error since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic. Fig. 3 shows the inference error for AoI values ${\delta=1,2,\ldots,50}$ and packet lengths $l=1,2,4,6,8,10$.
Using this AR process, we illustrate the performance of the optimal policy derived for the scheduling problem specified in (6)-(7) under time-invariant packet length selection.
Let the number of delay states be $N=2$, implying that the delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$. During the simulations, we assume $p_{12}=p_{21}$, ensuring that each delay state has a steady-state probability of $\frac{1}{2}$. The transmission delay $T_{i}(l)$ for each delay state is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ \lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (19)
where $\sigma$ is a scaling parameter. Additionally, the feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$. The maximum allowed packet length is $10$, and the buffer size is $B=75$.
We control the delay memory by varying the sum of transition probabilities $\alpha=p_{12}+p_{21}$. When $\alpha=1$, i.e., ${p_{12}=p_{21}=\frac{1}{2}}$, both delay states are equally likely in the $i$-th epoch, independent of the delay state in the $(i-1)$-th epoch. Consequently, the delay distribution is i.i.d.. However, as $\alpha$ deviates from $1$, the delay memory increases, and the delay state in the $(i-1)$-th epoch becomes more informative about the delay state in the $i$-th epoch.
@@
    Policy outlined in Theorem 1 with $l=1$.
@@
The result presented in Fig. 4 highlights that the optimal policy achieves a time-average inference error up to six-times lower than that of the generate-at-will, zero-wait, $l=1$ policy. Part of this improvement can be attained using a packet length of $l=1$ with the waiting times and buffer positions determined as described in Theorem 1. However, a substantial portion of the performance gain comes from selecting the appropriate packet length, accounting for the interplay between the transmission delay statistics and the packet length $l$.
@@
    Policy outlined in Theorem 1 with $l=5$. This policy accounts for the delay memory.
@@
    Policy outlined in [17, Theorem 1] with $l=5$. This policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
The result in Fig. (5) emphasizes the importance of accounting for delay memory when designing scheduling algorithms. The performance of both policies is quite similar when the delay state of the network is almost i.i.d., i.e., when $\alpha$ is around $1$. However, as $\alpha$ deviates from this value, the delay memory increases, and the advantage of the policy in Theorem 1 becomes apparent, with a significant performance gain of up to $11.6\%$.
@@ V-B Trace-Driven Evaluation
Figure 6: Structure of the cart-pole experiment.
@@
Ensuring the safe operation of robotic systems often requires evaluating their state based on observations from remote sources. This makes the prediction problem critical for monitoring the safety of such systems. We consider a specific scenario, illustrated in Fig. 6, where a camera remotely monitors a control system, e.g., cart-pole. The camera captures video frames and transmits them to a receiver over a network with distinct delay states. Due to the high dimensionality of the video data, transmitting the sequence of frames takes multiple time slots. As a result, the most recently delivered sequence of video frames is generated $\delta$ time slots ago. Upon reception, the frames undergo data pre-processing to extract the cart position and the pole angle information, ${X_{t-\delta}=\big{(}(\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta-l+1},Z_{t-\delta-l+1})\big{)}}$. Subsequently, a predictor uses the most recent available information to estimate the current pole angle and cart position ${Y_{t}=(\psi_{t},Z_{t})}$.
Figure 7: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and packet lengths ${l=1,2,4,6,8,10}$.
We consider the OpenAI CartPole-v1 task [47], where a Deep Q-Network (DQN) reinforcement learning algorithm [48] is used to control the force applied to a cart to prevent the attached pole from falling over. A pre-trained DQN neural network from [6] is employed for this control task. A time-series dataset containing the pole angle $\psi_{t}$ and cart position $Z_{t}$ is obtained from [6]. This dataset was generated by simulating 10,000 episodes of the OpenAI CartPole-v1 environment using the DQN controller.
@@
The predictor in this experiment is a Long Short-Term Memory (LSTM) neural network consisting of one input layer, one hidden layer with 64 LSTM cells, and a fully connected output layer. The first 72% of the dataset is used for training, and the remaining data is used for inference. The performance of the trained LSTM predictor is evaluated on the inference dataset. The inference error is shown in Fig. 7 for AoI values $\delta=1,2,\ldots,80$ and packet lengths $l=1,2,4,6,8,10$. We observe that the inference error is not a monotonic function of AoI $\delta$ for a given packet length $l$. Moreover, the inference error decreases as the packet length $l$ increases for a fixed AoI value $\delta$.
@@
Using this cart-pole experiment, we illustrate the performance of the optimal policy in Theorem 2, which is derived for the scheduling problem (14) under time-variable packet length selection.
We consider two delay states $N=2$. The delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$. We set $p_{12}=p_{21}$ to ensure equal steady-state probabilities for both delay states. The transmission delay $T_{i}(l)$ for each delay state is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (20)
@@
where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is identical to that in the AR process experiment. We set the sum of transition probabilities as $\alpha=p_{12}+p_{21}=\frac{1}{20}$.
@@ VI Conclusion
This paper studies a goal-oriented communication design problem for remote inference, where an intelligent neural network model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location. We derive two optimal scheduling policies under time-invariant and time-variable packet length selection. These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory. Finally, through the simulations, we demonstrate that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
@@ Appendix B Proof of Theorem 1
By using the linearity of expectation, we can obtain